OpenStudy (lillylilly1126):

Write an equation in slope intercept form of the line that passes through the given points. 23. (10,2) and (-2,-2)

OpenStudy (trojanpoem):

Slope - intercept form: \[y = mx + b\] Slope formula: \[\frac{ \Delta y }{ \Delta x } = \frac{ y_{2} - y_{1} }{ x_{2} - x_{1} }\] \[\frac{ \Delta y }{ \Delta x } = m = slope.\] \[slope = \frac{ -2 -2 }{ -2 -10 } = \frac{ -4 }{-12 } = \frac{ 1 }{ 3 }\] \[y = \frac{ 1 }{ 3}x + b\] From your points ( 10,2) , (-2,-2)\[-2 = \frac{ -2 }{ 3 } + b\] \[b = -2 +\frac{ 2 }{ 3 } = \frac{ -4 }{ 3 }\] \[y = \frac{ 1 }{ 3 }x -\frac{ 4 }{ 3 }\]